OpenStudy (anonymous):

Speacialty t-shirts are being sold online for $30 each plus a one-time handling fee of $2.75. The total cost is a function of the number of t-shirts bought.What function rule models the cost of the t-shirts? Evaluate the function of 5 t-shirts. 1.) 30t+2.75;$43.75 2.) 2.75t+30;$43.75 3.) 2.75t+30;$152.75 4.) 30t+2.75;$152.75 please help! i believe it is 3

OpenStudy (texaschic101):

30t + 2.75 5 tee shirts would be : 30(5) + 2.75 = 150 + 2.75 = 152.75

OpenStudy (texaschic101):

it is 4...because the tee shirts cost $5 a piece....plus the one time fee of $2.75

OpenStudy (texaschic101):

oops...the tee shirts cost $30 and for 5 shirts it is 150 plus the ine time fee of $2.75

OpenStudy (anonymous):

so it is1?

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

no..it is 4

OpenStudy (anonymous):

and how did you get this

OpenStudy (texaschic101):

ok....let me explain t = # of shirts It costs $30 per shirt plus a one time fee of $2.75 so your equation would be : 30t + 2.75 and if t = 5 (for 5 tee shirts) sub 5 in for t 30t + 2.75 = 30(5) + 2.75 = 150 + 2.75 = 152.75 so you have your equation and you have your total cost when you buy 5 tee shirts any better understanding ?